### v1

# Dynamic Obstacle Avoidance of Unmanned Surface Vehicles in Maritime Environments Using Gaussian Processes Based Motion Planning

Abstract
During recent years, unmanned surface vehicles are extensively utilised in a variety of maritime applications such as the exploration of unknown areas, autonomous transportation, offshore patrol and others. In such maritime applications, unmanned surface vehicles executing relevant missions that might collide with potential static obstacles such as islands and reefs and dynamic obstacles such as other moving unmanned surface vehicles. To successfully accomplish these missions, motion planning algorithms that can generate smooth and collision-free trajectories to avoid both these static and dynamic obstacles in an efficient manner are essential. In this article, we propose a novel motion planning algorithm named the Dynamic Gaussian process motion planner 2, which successfully extends the application scope of the Gaussian process motion planner 2 into complex and dynamic environments with both static and dynamic obstacles. First, we introduce an approach to generate safe areas for dynamic obstacles using modified multivariate Gaussian distributions. Second, we introduce an approach to integrate real-time status information of dynamic obstacles into the modified multivariate Gaussian distributions. Therefore, the multivariate Gaussian distributions with real-time statuses of dynamic obstacles can be innovatively added into the optimisation process of factor graph to generate an optimised trajectory. The proposed Dynamic Gaussian process motion planner 2 algorithm has been validated in a series of benchmark simulations in the Matrix laboratory and a dynamic obstacle avoidance mission in a high-fidelity maritime environment in the Robotic operating system to demonstrate its functionality and practicability.
keywords:
Unmanned surface vehicles , Gaussian process motion planner 2 , Multivariate Gaussian distribution , Dynamic obstacles avoidance.
1 Introduction
Benefit from the flexibility and versatility of USVs, they have been widely applied in a series of autonomous maritime missions in the past decades. The employment of USVs in autonomous maritime missions involves three fundamental components:
  • 1.
    First, sensing the surrounding maritime environments through onboard sensors.
  • 2.
    Second, planning safe and smooth trajectories based on sensed information.
  • 3.
    Third, navigating the platforms to follow the generated trajectories to perform relevant missions.
Sensing and navigation are effectively connected through planning, which is responsible for processing sensed information as well as for generating optimised trajectories for navigation. It is important to factor in not just static obstacles such as islands and reefs, but also dynamic obstacles such as other moving USVs during the motion planning processes of USVs. In addition, the motion planning algorithms must plan safe and smooth trajectories within an extreme short time after detecting any dynamic obstacle in maritime environments. Otherwise, USVs may have collided with dynamic obstacles despite that safe and smooth trajectories can be generated afterwards.
To our knowledge, the current mainstream motion planning algorithms are originally designed to generate safe trajectories between start and goal positions while avoiding static obstacles. The fundamental versions of the current mainstream motion planning algorithms do not take the real-time statuses of dynamic obstacles into account when planning trajectories. The current mainstream motion planning algorithms contain: 1) Gird-based algorithms, 2) Sampling-based algorithm, 3) Potential field algorithms, 4) Intelligent algorithms, 5) GP-based algorithms and others. Below is a summary of the current mainstream motion planning algorithms:
  • 1.
    The grid-based algorithms might require a relatively long computation time to find a feasible path due to the nature of the algorithms, especially when the dimension of the motion planning problems increases [1, 2, 3, 4, 5].
  • 2.
    The sampling-based algorithms cannot always guarantee the quality of the generated path. It is possible that the length of the generated path is relatively long. The smoothness of the generated path is also uncertain. Some sampling-based algorithms might require 1) a relatively long time to ensure the convergence because of their randomness and 2) some other time for post-processing path smoothing procedure [1, 2, 6, 7, 8, 9, 10].
  • 3.
    The potential field algorithms are normally limited by the local minima problems. Because of the nature of potential field algorithms, it is impossible to overcome the local minima problem without applying additional strategies [1, 2, 11, 12].
  • 4.
    The intelligent algorithms sometimes require a relatively long computation time and are limited by the local minima problems in some cases [1, 2, 13, 14, 15].
  • 5.
    The GP-based algorithms have several benefits among the mainstream motion planning algorithms [16, 17, 18, 19]. They can generate an optimised path in an efficient manner by smoothing and shortening the path during the planning process. Furthermore, they can eliminate the drawback of local minima by using the probabilistic inference [1, 19].
Table 1: A list of abbreviations in this article.
Abbreviations Definitions
2-D 2 dimensional
3-D 3 dimensional
k-D k dimensional
AIS Automatic identification system
AUV Autonomous underwater vehicle
CV Constant-velocity
CV-SL Constant-velocity straight-line
D-GPMP2 Dynamic Gaussian process motion planner 2
GPS Global positioning system
GP Gaussian process
GP-based Gaussian-process-based
GPMP2 Gaussian process motion planner 2
LTV-SDE Linear time-varying stochastic differential equation
MAP Maximum a posterior
MATLAB Matrix laboratory
ROS Robotics operating system
USV Unmanned surface vehicle
WAM-V 20 Wave adaptive modular vessel 20
Almost all the aforementioned mainstream motion planning algorithms have been improved and can avoid dynamic obstacles in some way. A series of examples is provided as follows:
  • 1.
    [20] has proposed an approach to improve the fundamental version of the grid-based motion planning algorithms and enable it to work in complex and dynamic environments.
  • 2.
    [21] has proposed an approach to integrate the real-time status information of dynamic obstacles into a potential field algorithm.
  • 3.
    [22] has proposed an approach to combine a sampling-based motion planning algorithm with the international regulations for preventing collisions at sea to avoid dynamic obstacles.
  • 4.
    [23, 24] have proposed a series of strategies for the sampling-based motion planning algorithms to avoid dynamic obstacles through 1) abandoning the minority of tree branches invalid by dynamic obstacles, 2) maintaining the rest majority of tree branches and 3) generating new collision-free tree branches.
  • 5.
    [25] has proposed a two-layer approach that combines an intelligent algorithm and a potential field algorithm to avoid dynamic obstacles.
  • 6.
    [26] has investigated how to use the GP-based algorithms to avoid dynamic obstacles and proposed a new factor graph structure.
Nevertheless, it still remains to be researched how to use the GP-based motion planning algorithms to overcome static and dynamic obstacles simultaneously. To bridge this research gap, this article has specifically focused on developing a novel motion planning algorithm for USVs to avoid both static and dynamic obstacles in maritime environments. The contributions of this work can be summarised as follows:
  • 1.
    First, an approach is introduced for generating a safe area for a moving USV. More specifically, this is achieved by incorporating the USV dimension specification from relevant sensor data into a multivariate Gaussian distribution’s variance matrix.
  • 2.
    Second, an approach is introduced for integrating various real-time status information of the moving USV into the established safe area. To be more precise, we introduce an approach for integrating the USV speed into the established safe area’s variance matrix. In addition, we design a function to enable both the rotation and translation of the established safe area based upon image processing techniques. Therefore we are able to align it with the course and position of the moving USV.
  • 3.
    Third, a new structure of the used optimisation tool is introduced to deal with the GP-based motion planning problem with both static and dynamic obstacles.
  • 4.
    Fourth, the proposed motion planning algorithm is thoroughly tested and validated in 1) a series of self-constructed simulation environments in MATLAB and 2) a high-fidelity maritime environment in ROS.
In the self-constructed simulation environments in MATLAB, we have tested and validated the performance of the proposed motion planning algorithm in a series of benchmark simulations including:
  • 1.
    A benchmark with different weights of dynamic obstacle likelihood.
  • 2.
    A benchmark with a dynamic obstacle with different velocities.
  • 3.
    A benchmark with dynamic obstacles with different areas.
  • 4.
    A benchmark with multiple dynamic obstacles.
  • 5.
    A benchmark with static and dynamic obstacles.
The results demonstrate the following:
  • 1.
    The proposed motion planning algorithm can generate a safe and smooth trajectory to avoid dynamic obstacles by considering their areas, directions, speeds and positions.
  • 2.
    The proposed motion planning algorithm can generate an optimised trajectory in an extremely short execution time.
  • 3.
    The proposed motion planning algorithm can be utilised in a variety of complex environments with multiple obstacles and different statuses.
Figure 1: Scenario demonstration in the Gazebo simulation environment in ROS: two USVs are approaching each other and at least one of them requires to take action to avoid the other. The high-fidelity maritime environment is provided by [27].
We have also tested and validated the proposed motion planning algorithm in a high-fidelity maritime environment in ROS. In ROS, augmented practicability is achieved based on the consideration of a series of physical properties such as hydrodynamic force, gravity and buoyancy [1, 27]. In comparison with the expensive testing process using a full-sized physical USV, validating the developed motion planning algorithm in ROS can effectively reduce the testing cost and maintain reliable test results simultaneously [1, 2]. A scenario demonstration in the Gazebo simulation environment in ROS is provided in Fig. 1.
The reminder of this article is organised as follows. Section 2 presents the mathematical model of the GPMP2 algorithm in detail. Section 3 presents the proposed D-GPMP2 algorithm in detail. Section 4 presents a new structure of the optimisation tool used in the research. Section 5 demonstrates a series of simulation results of the proposed motion planning algorithm in MATLAB. Section 6 demonstrates the practical performance of the proposed motion planning algorithm in ROS, while Section 7 concludes this article and proposes some future research directions.
2 GPMP2
This section presents the traditional GP-bsaed motion planning algorithm (GPMP2 algorithm) and summarises the previous work in [1, 2, 16, 17, 18, 19]. In general, the GPMP2 algorithm considers a motion planning problem as a standard optimisation problem. To be more specific, the GPMP2 algorithm views the potential sample trajectories as a GP prior, a conditional constraint that facilitates the capability of avoiding static obstacles as a static obstacle likelihood and the optimisation process as a MAP estimation to obtain an optimal solution [1, 19].
2.1 GP prior
The GP prior can generate a large amount of continuous-time sample trajectories, which has a high possibility of containing one or multiple trajectories that satisfied the predefined requirements. Continuous-time trajectories are assumed to be sampled from a continuous-time GP [16, 17, 19]:
$\theta(t)\sim\mathcal{GP}(\mu(t),K(t,t^{\prime})),$ (1)
where $\mu(t)$ is a vector-valued mean function $\mu=[\mu_{1}...\mu_{N}]^{T}$ and $K(t,t^{\prime})$ is a matrix-valued covariance function $K=[K(t_{i},t_{j})]|_{(0\leq i,j\leq N)}$.
In the GPMP2 algorithm, this continuous-time GP is considered to be generated by a LTV-SDE. To be more specific, the following equation is used to describe the LTV-SDE:
$\dot{\theta}(t)=A(t)\theta(t)+u(t)+F(t)w(t),$ (2)
where $u(t)$ is a known exogenous input, $A(t)$ and $F(t)$ are system matrices and $w(t)$ represents the white process noises. The white process noises is a GP with zero mean:
$w(t)\sim GP(\textit{0},Q_{c}\delta(t-t^{\prime})),$ (3)
where $Q_{c}$ is a positive semi-definite time-varying power-spectral density matrix and $\delta(t-t^{\prime})$ is the Dirac delta function.
The solution of the continuous-time GP can be obtained by the process described in [19, 28] as:
$u(t)=\Phi(t,t_{0})u_{0}+\int_{t_{0}}^{t}\Phi(t,s)u(s)\textit{ds}$ (4)
$\begin{gathered}K(t,t^{\prime})=\Phi(t,t_{0})K_{0}\Phi(t^{\prime},t_{0})^{T}+\\ \int_{t_{0}}^{min(t,t^{\prime})}\Phi(t,s)F(s)Q_{c}F(s)^{T}\Phi(t,s)^{T}\textit{ds},\end{gathered}$ (5)
where $u_{0}$ and $K_{0}$ are initial mean and covariance of the first state, respectively. $\Phi(t,s)$ is the state transition matrix from time t to time s.
As a result, the GP prior distribution can be expressed by the following equation:
$p(\theta)\propto exp\{-\frac{1}{2}||\theta-u||^{2}_{K}\}.$ (6)
Sample trajectories from the GP prior provide the user with a large number of trajectory options from which to choose the optimised one depending on their requirements specified in the likelihoods.
2.2 Static obstacle likelihood
In the GPMP2 algorithm, a static obstacle likelihood that facilitates the event of avoiding static obstacles $l(\theta;\textit{$e_{st\_obs}$})$ is used and can be defined as a distribution in the exponential family as below:
$l(\theta;\textit{$e_{st\_obs}$})=exp\{-\frac{1}{2}||h_{st\_obs}(\theta)||^{2}_{\sigma_{st\_obs}}\},$ (7)
where $e_{st\_obs}$ represents the event of avoiding static obstacles, the definition of matrix $\sigma_{\textit{st\_obs}}$ can be found in [17, 19] and $h_{st\_obs}(\theta)$ is a vector-valued cost function:
$h_{st\_obs}(\theta_{i})=[c(d(\theta_{i}))],$ (8)
where $c$ is the hinge loss function defined by the equation:
$c(d)=\left\{\begin{array}[]{cr}-d+\epsilon,&d\leq\epsilon\\ 0,&d>\epsilon\\ \end{array}\right.,$ (9)
where $d$ is the signed distance of a point and $\epsilon$ is the safety distance. In general, the safety distance $\epsilon$ is defined based on user experience and the signed distance function $d$ is stated as:
$d=D-\overline{D},$ (10)
where $D$ is the Euclidean distance transforms function (also named as distance field). $\overline{D}$ is the complement of distance field $D$.
Based on the specific requirements for achieving different purposes, GPMP2 can incorporate various likelihoods. In this article, Section 3 combines the GPMP2 algorithm with a dynamic obstacle likelihood to allow it to avoid both static and dynamic obstacles.
2.3 MAP estimation
The MAP estimation is an estimate of an unknown quantity, which equals the mode of the posterior distribution [29]. Now it is possible to perform a Bayesian inference to obtain an optimised trajectory using the knowledge provided in Sections 2.1 and 2.2: 1) a GP prior that includes a variety of sample trajectories from a start point to a goal point and 2) a likelihood that facilitates the event of avoiding static obstacles. Thus, the Bayesian inference of this problem can be defined as:
$\begin{gathered}\theta_{0:N}^{*}=\mathop{\arg\max}_{\theta}p(\theta|{e})=\mathop{\arg\max}_{\theta}\;p(\theta)p(e_{st\_obs}|\theta).\end{gathered}$ (11)
Figure 2: Flow diagram of the proposed D-GPMP2 algorithm. The innovative part enhances the original GPMP2 algorithm into the proposed D-GPMP2 algorithm that can avoid static and dynamic obstacles simultaneously.
The MAP estimation of a GP can be converted into a least squares problem as described in [19], which has been well-studied for many years and there are many methods to solve least squares problem including but not limited to Gaussian-Newton method and Levenberg–Marquardt method [30, 31]. In this work, Levenberg–Marquardt method is used to solve the least squares problem in any GP-based motion planning algorithm.
3 D-GPMP2
This section presents the proposed D-GPMP2 algorithm, in detail, which includes two new parts added into the original GPMP2 algorithm: 1) Dynamic obstacle area modelling part and 2) Dynamic obstacle velocity modelling part.
Figure 3: (a) USV dimension specification from AIS data. (b) USV safe area based on the multivariate Gaussian distribution radius $R^{\prime}$ and $R^{\prime}$ is proportional to the self-defined safe radius $R$. The USV model in the above sub-figures is from [27].
For the dynamic obstacle area modelling, we propose a method that combines the AIS data and multivariate Gaussian distribution to construct a safe area for the target USV.
For the dynamic obstacle velocity modelling, we propose: 1) a method which can integrate the speed of the target USV into the multivariate Gaussian distribution and 2) a function that can rotate the multivariate Gaussian distribution with speed information to the correct direction and translate the rotated multivariate Gaussian distribution to the correct location.
Based on these two new parts, a dynamic obstacle likelihood can be constructed and the MAP estimation of the proposed D-GPMP2 algorithm considers both the static obstacle likelihood and the dynamic obstacle likelihood. By adding dynamic obstacle likelihood, the proposed D-GPMP2 algorithm gives the original GPMP2 algorithm the capability of avoiding static and dynamic obstacles. The flow diagram of the proposed D-GPMP2 algorithm is illustrated in Fig. 2.
3.1 Dynamic obstacle area modelling
The AIS is a fundamental equipment mounted on the USVs to display maritime traffic information and communicate with other USVs. In general, maritime traffic information information from the AIS contains: 1) the target USV position from the GPS, 2) the target USV dimension, 3) the target USV speed, 4) the target USV course (also known as target USV direction) and others [32].
In maritime environments, other moving USVs are the most common dynamic obstacles for a USV. Fig. 3 (a) illustrates the information from the AIS that describes the dimension specification of a target USV: $A$ represents the distance from forward perpendicular, $B$ represents the distance from after perpendicular, $C$ represents the distance inboard from port side and $D$ represents the distance inboard from starboard side. In the meantime, Fig. 3 (b) demonstrates the USV safe area determined by a self-defined safe radius $R$ of the USV using the equation below:
$R=\zeta\cdot\frac{\overbrace{A+B}^{length}+\overbrace{C+D}^{width}}{2},$ (12)
where $\zeta$ is an adjustable positive scaling factor, $[A+B]$ represents the length of the USV, $[C+D]$ represents the width of the USV and the USV safe area is constructed by a multivariate Gaussian distribution. Furthermore, the self-defined safe radius $R$ should ensure the multivariate Gaussian distribution at least fully covers the USV.
Figure 4: (a) The safe area of the Heron USV in MATLAB from the 2-D perspective. (b) The safe area of the Heron USV in MATLAB from the 3-D perspective. (c) The safe area of the WAM-V 20 USV in MATLAB from the 2-D perspective. (d) The safe area of the WAM-V 20 USV in MATLAB from the 3-D perspective. In the colour bars, a larger value indicates a higher risk and a smaller value indicates a lower risk.
Table 2: USV dimension specification [2, 33].
USV Name USV Length USV Width
WAM-V 20 USV 6 [m] 3 [m]
Heron USV 1.35 [m] 0.98 [m]
Generally, the k-D Gaussian distribution can be written as [34]:
$X\sim\mathcal{N}_{k}(\mu,\Sigma),$ (13)
where $\mu$ is the mean vector and $\Sigma$ is the covariance matrix. The probability density function of the k-D Gaussian distribution in (13) can be expressed as:
$\small f_{k}(X_{1},...,X_{k})=\frac{1}{\sqrt{(2\pi)^{k}|\Sigma|}}\exp[-\frac{1}{2}(X-\mu)^{T}\Sigma^{-1}(X-\mu)].$ (14)
Due to the motion planning of USVs lies in the 2-D environments, therefore the 2-D Gaussian distribution is employed. In the 2-D situation, the probability density function can be transformed into the following form:
$\begin{split}f_{2}(X,Y)=\frac{1}{2\pi\sigma_{X}\sigma_{Y}\sqrt{1-\rho^{2}}}\exp\{-\frac{1}{2(1-\rho^{2})}\\ [\frac{(X-\mu_{X})^{2}}{\sigma_{X}^{2}}+\frac{(Y-\mu_{Y})^{2}}{\sigma_{Y}^{2}}-\frac{2\rho(X-\mu_{X})(Y-\mu_{Y})}{\sigma_{X}\sigma_{Y}}]\},\end{split}$ (15)
Figure 5: A comparison of the models constructed by (19) and (20). The sub-figures (a) - (e) demonstrate the models constructed by (19). The sub-figures (f) - (j) demonstrate the models constructed by (20). In both comparative groups, the speed of the target USV changes from 0 [m/s] to 15 [m/s]. In the colour bars, a larger value indicates a higher risk and a smaller value indicates a lower risk.
Figure 6: (a) - (c) demonstrates of the working process of $f_{crs\_pos}$ rotates an image matrix. (d) - (f) demonstrates of the working process of $f_{crs\_pos}$ translates an image matrix. The risk at the point close to the dark red region is relatively high; on the other hand, the risk at the point close to the dark blue region is relatively low.
Figure 7: Demonstration of the modelling results of target USVs with different velocities and positions in MATLAB from the 2-D perspective and the 3-D perspective. In the colour bars, a larger value indicates a higher risk and a smaller value indicates a lower risk.
where $\rho$ is the correlation coefficient of $X$ and $Y$. The mean vector $\mu$ can determine the position of the 2-D Gaussian distribution and the covariance matrix $\Sigma$ can influence the shape of the 2-D Gaussian distribution:
$\mu=\begin{bmatrix}\mu_{X}\\ \mu_{Y}\end{bmatrix},\Sigma=\begin{bmatrix}\sigma_{X}^{2}&\rho\sigma_{X}\sigma_{Y}\\ \rho\sigma_{X}\sigma_{Y}&\sigma_{Y}^{2}\end{bmatrix},$ (16)
where $\rho$ is the correlation coefficient between $X$ and $Y$, $\mu_{X}$ represents the $X$ position of the 2-D Gaussian that equals to $X$ position of the GPS data $x_{gps}$, $\mu_{Y}$ represents the $Y$ position of the 2-D Gaussian that equals to the $Y$ position of the GPS data $y_{gps}$, $\sigma_{X}^{2}$ represents the extension of 2-D Gaussian along $X$ axis and $\sigma_{Y}^{2}$ represents the extension of 2-D Gaussian along $Y$ axis [35].
As a result, the relationships between $\sigma_{X}^{2}$, $\sigma_{Y}^{2}$ and $R$ can be described as below:
$\sigma_{X}^{2}=\sigma_{Y}^{2}=R^{2},$ (17)
Further, the relationships between $\sigma_{X}^{2}$, $\sigma_{Y}^{2}$ and AIS data that describes the dimension specification of the target USV can be transformed as below:
$\sigma_{X}^{2}=\sigma_{Y}^{2}=[\frac{\zeta\cdot(A+B+C+D)}{2}]^{2}.$ (18)
When constructing the safe area of a USV, the USV is always placed at the centre of the image and will be translated to the correct position according to GPS data later. To provide a more intuitive understanding of the proposed dynamic obstacle area modelling method, the examples in Fig. 4 demonstrates the modelling results of the safe areas of USVs with different dimensions. Meanwhile, Table 2 provides the dimension specifications of the modelled USVs.
3.2 Dynamic obstacle velocity modelling
To reduce the risk of colliding with a target USV, it is necessary to integrate the position and velocity information of the target USV into the model constructed in Section 3.1. To achieve this goal, we need to use the target USV position from the GPS data and target USV speed and course from the AIS data. By incorporating the information into a new model, the proposed D-GPMP2 algorithm can use the new model to generate a trajectory to avoid any potential movement of the target USV.
The speed of a target USV from the AIS data can be integrated into the dynamic obstacle area model by using the following equation:
$\begin{bmatrix}\sigma_{X}^{2}\\ \sigma_{Y}^{2}\end{bmatrix}=\begin{bmatrix}R^{2}\\ (\eta\cdot v+1)\cdot R^{2}\end{bmatrix},$ (19)
where $\eta$ is a adjustable positive scaling factor and $v$ is the target USV speed from the AIS data. Now the safe area constructed by the method described in Section 3.1 is extended along the positive half axis of $Y$ axis as a result of the influence of (19). Nevertheless, we notice that the increased length in each meter per second of the model constructed by (19) gradually decreases with the speed gradually increasing. Therefore, we add a self-defined exponential function into (19) to calibrate this error:
$\begin{bmatrix}\sigma_{X}^{2}\\ \sigma_{Y}^{2}\end{bmatrix}=\begin{bmatrix}R^{2}\\ (\eta\cdot v+1)\cdot a^{(b\cdot v)}\cdot R^{2}\end{bmatrix}$ (20)
As demonstrated in the first comparative group in Fig. 5 (a) - (e), the models constructed by (19) no longer change significantly with the speed increases. Nevertheless, as demonstrated in the second comparative group in Fig. 5 (f) - (j), the models constructed by (20) with the self-defined exponential function reduce the error. It is worth noting that the model constructed by our method is extended both the forward and backward directions of velocity. Due to the possibility of small errors in GPS positioning, the model expands in the backward direction of velocity to ensure the obstacle can be avoided from behind safely.
Next, the model constructed by (20) needs to be rotated towards the target USV course and translated to the target USV position. To achieve these goals, a function is designed as below:
$M_{N\times N}=f_{crs\_pos}(M_{N\times N},\alpha,x_{gps},y_{gps}),$ (21)
where $M_{N\times N}$ is a $N$-by-$N$ image matrix that contains the model constructed by (20). More specifically, there are two main steps when the function in (21) rotates an image matrix: 1) rotating the original image around its centre with a specified angle $\alpha$, where $\alpha$ equals to the difference between the target USV course and the positive half axis of $Y$ axis and 2) interpolating zero values into the pixels that are inside the new boundary of the image matrix but outside the rotated image. Similarly, there are two main steps when the function designed in (21) translates an image matrix: 1) translating the centre of the original image to the target USV position [$x_{gps}$, $y_{gps}$] and 2) interpolating zero values into the pixels that are inside the new boundary of the image matrix but outside the translated image.
Fig. 6 provides a series of examples to give a more intuitive understanding of the working process of $f_{crs\_pos}$, while Fig. 7 demonstrates the modelling results of target USVs with different velocities and positions based on (20) and (21).
Figure 8: (a) Object modelled by the artificial potential field. (b) Object modelled by the 2-D Gaussian. In both sub-figures, a larger value indicates a higher risk and a smaller value indicates a lower risk.
3.3 A comparison between multivariate Gaussian and artificial potential field
It was found that the artificial potential field algorithm has been used as a mainstream method of building object models in many previous studies deal with dynamic object avoidance in motion planning and a famous example can be found in [21]. In order to prove the effectiveness of the proposed method and a comparison of an object modelled by using the artificial potential field and 2-D Gaussian is provided in Fig. 8. It is worth noting that the 2-D Gaussian can generate a model with a clearer boundary and a clear boundary helps to generate a trajectory that can avoid dynamic objects in a more efficient manner.
3.4 Dynamic obstacle likelihood
Based on the aforementioned information, a dynamic obstacle likelihood that considers 1) the safe areas, 2) velocities and 3) positions of the target USVs can be written as a distribution in the exponential family as [1, 19]:
$l(\theta;\textit{$e_{dy\_obs}$})=exp\{-\frac{1}{2}||h_{dy\_obs}(\theta)||^{2}_{\sigma_{dy\_obs}}\},$ (22)
where $e_{dy\_obs}$ represents the event of avoiding dynamic obstacles, the definition of matrix $\sigma_{\textit{dy\_obs}}$ is similar with $\sigma_{\textit{st\_obs}}$ and $h_{dy\_obs}(\theta)$ is a vector-valued cost function:
$h_{dy\_obs}(\theta_{i})=[f_{crs\_pos}(f_{2}(\theta_{i}),\alpha,x_{gps},y_{gps}))],$ (23)
where the mean vector $\mu$ and the covariance matrix $\Sigma$ of the 2-D Gaussian probability density function $f_{2}$ are:
$\footnotesize\mu=\begin{bmatrix}x_{gps}\\ y_{gps}\end{bmatrix},$ (24)
$\footnotesize\Sigma=\begin{bmatrix}\frac{1}{4}\zeta^{2}(A+B+C+D)^{2}&0\\ 0&\frac{1}{4}\zeta^{2}(\eta v+1)a^{bv}(A+B+C+D)^{2}\end{bmatrix}.$ (25)
3.5 MAP estimation
Based on the static obstacle likelihood and the dynamic obstacle likelihood, the Bayesian inference of the proposed D-GPMP2 algorithm can then be defined as:
$\begin{gathered}\theta_{0:N}^{*}=\mathop{\arg\max}_{\theta}\;p(\theta)p(e_{st\_obs}|\theta)p(e_{dy\_obs}|\theta).\end{gathered}$ (26)
In the proposed D-GPMP2 algorithm, we also use the Levenberg–Marquardt method to solve the least squares problem [1, 19, 31].
Figure 9: Overall structure of the factor graph. $f^{p}_{0}(\theta_{0})$ and $f^{p}_{N}(\theta_{N})$ represent the regular prior factors, $f^{gp}_{i}(\theta_{i},\theta_{i+1})$ represent the GP prior factors, $f^{st\_obs}_{i}(\theta_{i})$ represent the static obstacle likelihood factors, $f^{st\_obs}_{\tau_{j}}(\theta_{i},\theta_{i+1})$ represent the interpolated static obstacle factors, $f^{dy\_obs}_{i}(\theta_{i})$ represent the dynamic obstacle likelihood factors and $f^{dy\_obs}_{\tau_{j}}(\theta_{i},\theta_{i+1})$ represent the interpolated dynamic obstacle factors.
4 Factor graph
In the proposed D-GPMP2, factor graph is used as a optimisation tool to deal with the MAP estimation and a new structure of it is designed to calculate the optimised trajectory to avoid static and dynamic obstacles simultaneously. More specifically, this new structure includes three categories of factors [1, 19]:
  • 1.
    Prior factors such as 1) the regular prior factors and 2) the GP prior factors.
  • 2.
    Static obstacle likelihood factors such as 1) the regular static obstacle likelihood factors and 2) the interpolated static obstacle likelihood factors.
  • 3.
    Dynamic obstacle likelihood factors such as 1) the regular dynamic obstacle likelihood factors and 2) the interpolated dynamic obstacle likelihood factors.
Fig. 9 demonstrates the overall structure of the factor graph used in this article. Generally, a factor graph $G$ is a bipartite graph and can be defined as:
$G=\{\Theta,\mathcal{F},\mathcal{E}\},$ (27)
where $\Theta$ are a set of variable nodes, $\mathcal{F}$ are factor nodes attached to the corresponding variable nodes via $\mathcal{E}$ that are edges that connect the variable nodes and factor nodes. The factorisation of the posterior in our problem can be formulated as:
$p(\theta|e_{s},e_{d})\propto\prod^{M}_{m=1}f_{m}(\Theta_{m}),$ (28)
where $f_{m}$ are factors on variable subset $\Theta_{m}$ [1, 19].
There are static obstacle likelihood factors in our motion planning problem regardless of whether there are static obstacles. And if there are no static obstacles, a signed distance field is generated based on an empty binary map.
5 Simulations and discussions
This section demonstrates the performance of the proposed motion planning algorithm on dynamic obstacle avoidance under different situations.
Table 3: Specification of the used hardware platform.
Device Description Quantity
Processor 2.3-GHz Intel Core i7-11800H 16
RAM 8 GB 1
5.1 Simulation details
In general, the proposed motion planning algorithm has been tested using five different categories of benchmark simulations:
  • 1.
    The first benchmark simulation validated the performance of the proposed motion planning algorithm when different weights of dynamic obstacle likelihood were chosen.
  • 2.
    The second benchmark simulation validated the performance of the proposed motion planning algorithm when the velocity of the dynamic obstacle was adjusted. More specifically, two quantitative sub-simulations were conducted with only the speed or direction of the dynamic obstacle adjusted.
  • 3.
    The third benchmark simulation validated the performance of the proposed motion planning algorithm when the area of the dynamic obstacle was adjusted.
  • 4.
    The fourth benchmark simulation validated the performance of the proposed motion planning algorithm in different multiple dynamic obstacles situations.
  • 5.
    The fifth benchmark simulation validated the performance of the proposed motion planning algorithm when both static and dynamic obstacles required to be avoided in complex maritime environments.
In all the benchmark simulations, the proposed motion planning algorithm was initialised with a CV-SL trajectory based on the system dynamics model described in Section 5.2. Table 3 demonstrates the specification of the used hardware platform for running the benchmark simulations, while Fig. 10 demonstrates the maritime map used in the fifth benchmark simulation. Meanwhile, the obstacle angle in the following benchmark simulations is the target USV course.
Figure 10: (a) demonstrates the geographical map of Roadford Lake, Devon, UK [36]. (b) demonstrates the binary map about the Lake generated by us. A square part of the binary map in (b) is used in the fifth benchmark simulation.
Figure 11: Demonstration of the simulation results with different weights of dynamic obstacle likelihood: From (a) to (b), the weight of dynamic obstacle likelihood is 0.5 and 0.005, respectively. The start position is represented in green, the goal position is represented in red and the trajectory is represented in purple. In the colour bars, a larger value indicates a higher risk and a smaller value indicates a lower risk.
5.2 System dynamics model
The dynamics of the robot platform is represented with the double integrator linear system with additional white noise on acceleration [1, 17, 19]. A CV prior is used because it can minimise acceleration along the trajectory, reduce energy consumption and increase the smoothness of the trajectory. Thus, the trajectory is generated by LTV-SED with parameters:
$A=\begin{bmatrix}0&I\\ 0&0\end{bmatrix},x(t)=\begin{bmatrix}r(t)\\ v(t)\end{bmatrix},F(t)=\begin{bmatrix}0\\ I\end{bmatrix},u(t)=0,$ (29)
and given $\Delta t_{i}=t_{i+1}-t_{i}$,
$\small\Phi(t,s)=\begin{bmatrix}I&(t-s)I\\ 0&I\end{bmatrix},Q_{i,i+1}=\begin{bmatrix}\frac{1}{3}\Delta t_{i}^{3}Q_{C}&\frac{1}{2}\Delta t_{i}^{2}Q_{C}\\ \frac{1}{2}\Delta t_{i}^{2}Q_{C}&\Delta t_{i}Q_{C}\end{bmatrix},$ (30)
The sample trajectories from the GP prior is centred around a zero-acceleration trajectory.
5.3 Simulation with different weights of dynamic obstacle likelihood
In this subsection, we demonstrate the influence of the weight of the dynamic obstacle likelihood in the proposed D-GPMP2 algorithm.
We utilised the proposed D-GPMP2 algorithm to solve a motion planning problem with a dynamic obstacle located at a point between the start position and goal position and moving towards the start position. The major parameters used in the proposed D-GPMP2 algorithm when solving this motion planning problem is demonstrated in Table 4. In this simulation, we gradually adjusted the weight of the dynamic obstacle likelihood from a relatively larger value to a relatively smaller value and observed the change of the generated trajectory. As can be seen in Fig. 11 (a) - (b), the capability of the proposed D-GPMP2 algorithm to avoid dynamic obstacle gradually improved when the weight of the dynamic obstacle likelihood gradually decreased.
To summarise, the performance of the proposed D-GPMP2 algorithm to avoid dynamic obstacle is adjustable. Further, this characteristic can help us choose different obstacle avoidance strategies in different situations. An example would be:
  • 1.
    We are able to enhance the capability of the proposed D-GPMP2 algorithm of avoiding dynamic obstacle when the obstacle is moving towards our platform with a relatively fast speed.
  • 2.
    We are also able to improve the travel efficiency by maintaining the generated path straight if the dynamic obstacle is actively avoiding our platform.
5.4 Simulation with a dynamic obstacle with different velocities
In this subsection, we demonstrate the performance of the proposed D-GPMP2 algorithm when adjusting the velocity of the dynamic obstacle and maintaining the other parameters constant. More specifically, two quantitative simulations were conducted in order to demonstrate the influence of the speed and direction of the velocity of the dynamic obstacle on the generated trajectory.
Table 4: List of parameters for the simulation with different weights of dynamic obstacle likelihood.
Number Parameter name Parameter value
1 Map size 100 $\times$ 100 [m2]
2 Start position [20, 20]
3 Goal position [80, 80]
4 Obstacle position [50, 50]
5 obstacle speed 2 [m/s]
6 Obstacle angle 225 [degree]
7 Obstacle length 6 [m]
8 Obstacle width 3 [m]
9 Safe radius 9 [m]
5.4.1 Simulation with a dynamic obstacle with different speeds
This simulation used the proposed D-GPMP2 algorithm to solve a motion planning problem with a dynamic obstacle located between the start position and goal position and moving towards the start position. Table 5 demonstrates the major parameters used in the proposed D-GPMP2 algorithm when solving this motion planning problem. While running different simulation trials, we gradually increased the speed of the dynamic obstacle from a relatively smaller value to a relatively larger value and observed the change on the generated trajectory. As can be seen in the simulation results from Fig. 12 (a) - (d), increasing the speed of the dynamic obstacle caused the generated trajectory to avoid the dynamic obstacle earlier.
Figure 12: Demonstration of the simulation results with a dynamic obstacle with different speeds: From (a) to (d), the velocity of the dynamic obstacle is 0 [m/s], 5 [m/s], 10 [m/s] and 15 [m/s], respectively. The start position is represented in green, the goal position is represented in red and the trajectory is represented in purple. In the colour bars, a larger value indicates a higher risk and a smaller value indicates a lower risk.
Figure 13: Demonstration of the simulation results with a dynamic obstacle with different directions: From (a) to (d), the angle of the dynamic obstacle is 210 [degree], 240 [degree], 300 [degree] and 330 [degree], respectively. The start position is represented in green, the goal position is represented in red and the trajectory is represented in purple. In the colour bars, a larger value indicates a higher risk and a smaller value indicates a lower risk.
Table 5: List of parameters for the simulation with a dynamic obstacle with different speeds.
Number Parameter name Parameter value
1 Map size 100 $\times$ 100 [m2]
2 Start position [10, 50]
3 Goal position [90, 50]
4 Obstacle position [50, 50]
5 Obstacle angle 270 [degree]
6 Obstacle length 6 [m]
7 Obstacle width 3 [m]
8 Safe radius 9 [m]
9 Dynamic obstacle likelihood weight 0.005
In summary, the proposed D-GPMP2 algorithm can automatically find a suitable moment to avoid dynamic obstacle according to the speed of the dynamic obstacle. This characteristic ensures the safety of the generated trajectory as well as improves the efficiency of the proposed D-GPMP2 algorithm. As a result, the proposed D-GPMP2 algorithm is adaptable for motion planning problems involving dynamic obstacles with various speeds.
5.4.2 Simulation with a dynamic obstacle with different directions
The proposed D-GPMP2 algorithm was used in this simulation to solve a motion planning problem with a dynamic obstacle passing by between the start position and goal position with different angles. The major parameters used in the proposed D-GPMP2 algorithm when solving this motion planning problem are demonstrated in Table 6. In this simulation, we gradually adjusted the angle of the dynamic obstacle when it was passing by between the start position and goal position and observed the change on the generated trajectory. As can be seen in the simulation results in Fig. 13 (a) - (d), the curved part of the generated trajectory is always located at a position in the opposite direction of the velocity of the dynamic obstacle.
In summary, the proposed D-GPMP2 algorithm can generate a trajectory to avoid a dynamic obstacle through bypassing it from behind at a suitable moment. This characteristic reduces the collision risk and improves the efficiency of the proposed D-GPMP2 algorithm. Consequently, the proposed motion planning algorithm is adaptable to problems involving dynamic obstacles moving in different directions.
Table 6: List of parameters for the simulation with dynamic obstacles with different directions.
Number Parameter name Parameter value
1 Map size 100 $\times$ 100 [m2]
2 Start position [50, 10]
3 Goal position [50, 90]
4 Obstacle position [50, 50]
5 Obstacle speed 5 [m/s]
6 Obstacle length 6 [m]
7 Obstacle width 3 [m]
8 Safe radius 9 [m]
9 Dynamic obstacle likelihood weight 0.005
Table 7: List of parameters for the simulation with dynamic obstacles with different areas.
Number Parameter name Parameter value
1 Map size 100 $\times$ 100 [m2]
2 Start position [50, 90]
3 Goal position [50, 10]
4 Obstacle position [50, 50]
5 Obstacle angle 180 [degree]
6 Obstacle speed 5 [m/s]
7 Dynamic obstacle likelihood weight 0.005
Figure 14: Demonstration of the simulation results with dynamic obstacles with different areas: (a) The length of the dynamic obstacle is 6 [m], the width of the dynamic obstacle is 3 [m] and the safe radius of the dynamic obstacle is 9 [m]. (b) The length of the dynamic obstacle is 9 [m], the width of the dynamic obstacle is 6 [m] and the safe radius of the dynamic obstacle is 15 [m]. The start position is represented in green, the goal position is represented in red and the trajectory is represented in purple. In the colour bars, a larger value indicates a higher risk and a smaller value indicates a lower risk.
Figure 15: Demonstration of the simulation results with multiple dynamic obstacles with different areas and velocities: (a) The position of the first dynamic obstacle is [40, 45] and the position of the second dynamic obstacle is [60, 55]. Furthermore, the angle of the two dynamic obstacles is 225 [degree]. (b) The position of the first dynamic obstacle is [40, 45] and the position of the second dynamic obstacle is [60, 55]. Furthermore, the angle of the two dynamic obstacles is 135 [degree] and 315 [degree], respectively. The start position is represented in green, the goal position is represented in red and the trajectory is represented in purple. In the colour bars, a larger value indicates a higher risk and a smaller value indicates a lower risk.
5.5 Simulation with dynamic obstacles with different areas
In this subsection, we investigate how the proposed D-GPMP2 algorithm performs if the dynamic obstacle area changes while the other parameters remain constant.
In this simulation, the proposed D-GPMP2 algorithm was applied to a motion planning problem with a dynamic obstacle between the start position and goal position and moving towards the start position. Table 7 demonstrates the major parameters used in the proposed D-GPMP2 algorithm when solving this motion planning problem. We observed the change on the generated trajectory after gradually increasing the area of the dynamic obstacle from a relatively smaller value to a relatively larger value during a series of simulation trials. The simulation results in Fig. 14 (a) - (b) demonstrate that an increase in the area of the dynamic obstacle results in an earlier avoidance of the obstacle on the generated trajectory. The reason for this is that a dynamic obstacle with a similar speed and a larger area is more likely to collide.
In summary, the D-GPMP2 algorithm can guarantee the safety and efficiency of obstacle avoidance by taking into account the area of dynamic obstacles.
5.6 Simulation with multiple dynamic obstacles
In this subsection, we demonstrate the performance of the proposed D-GPMP2 algorithm in handling multiple dynamic obstacles.
In this simulation, two motion planning problems involving two dynamic obstacles were resolved using the proposed D-GPMP2 algorithm:
  • 1.
    In the first motion planning problem, two dynamic obstacles move towards the start position continuously.
  • 2.
    In the second motion planning problem, two dynamic obstacles pass between the start position and goal position in opposite directions.
Table 8 demonstrates the major parameters used in the proposed D-GPMP2 algorithm in these motion planning problems. The simulation results in Fig. 12 (a) - (b) demonstrate that the generated trajectories can successfully avoid the dynamic obstacles in both the motion planning problems.
Figure 16: Demonstration of the simulation result with both static and dynamic obstacles. (a) The trajectory avoids the static obstacle. (b) The trajectory avoids the dynamic obstacle. The start position is represented in green, the goal position is represented in red and the trajectory is represented in yellow. In the colour bar of sub-figure (a), a larger value indicates a lower risk and a smaller value indicates a higher risk. In the colour bar of sub-figure (b), a larger value indicates a higher risk and a smaller value indicates a lower risk.
In summary, the proposed D-GPMP2 algorithm is adaptable not only to single dynamic obstacle situations, but also to multiple dynamic obstacle situations.
Table 8: List of parameters for the simulation with multiple dynamic obstacles.
Number Parameter name Parameter value
1 Map size 100 $\times$ 100 [m2]
2 Start position [20, 20]
3 Goal position [80, 80]
4 Obstacle speed 3 [m/s]
5 Obstacle length 1.5 [m]
6 Obstacle width 1 [m]
7 Safe radius 2.5 [m]
8 Dynamic obstacle likelihood weight 0.005
Table 9: List of parameters for the simulation with static and dynamic obstacles.
Number Parameter name Parameter value
1 Map size 500 $\times$ 500 [m2]
2 Start position [310, 130]
3 Goal position [210, 260]
4 Obstacle position [310, 153]
5 Obstacle speed 3 [m/s]
6 Obstacle length 15 [m]
7 Obstacle width 8 [m]
8 Safe radius 23 [m]
9 Obstacle angle 225 [degree]
10 Dynamic obstacle likelihood weight 0.003
11 Static obstacle likelihood weight 0.05
12 Safe distance 20 [m]
Table 10: Summary of the simulation result with static and dynamic obstacles
Number Measurement Value
1 Path length 201.0 [m]
2 Static obstacle likelihood time cost 3.9 [ms]
3 Dynamic obstacle likelihood time cost 9.7 [ms]
4 Factor graph time cost 13.9 [ms]
5 MAP estimation time cost 11.8 [ms]
5.7 Simulation with static and dynamic obstacles.
This subsection describes the performance of the proposed D-GPMP2 algorithm when dealing with static and dynamic obstacles together in a complex and dynamic maritime environment.
Figure 17: Demonstration of the selected USV platform in the high-fidelity maritime environment in ROS Gazebo [27].
Figure 18: The connection between MATLAB (the proposed D-GPMP2 algorithm) and ROS Gazebo (WAM-V 20 USVs).
Figure 19: The storyboards of the dynamic obstacle avoidance mission in a high-fidelity maritime environment in ROS Gazebo using the proposed D-GPMP2 algorithm. The start and goal positions of our USV are [50, 10] and [50, 90], respectively. The initial position of the dynamic obstacle is [50, 50]. From (a) to (l), the sub-figures demonstrate the positions of our USV and the dynamic obstacle when time equals to 0 [s], 2 [s], 4 [s], 6 [s], 8 [s], 10 [s], 12 [s], 14 [s], 16 [s], 18 [s], 20 [s] and 22 [s] from the third-person perspective. The high-fidelity maritime environment is provided by [27].
This simulation used the proposed D-GPMP2 algorithm to solve a motion planning problem that involves dynamic obstacles passing between the start position and goal position as well as static obstacles such as islands surrounding the area. Table 9 demonstrates the major parameters used in the proposed D-GPMP2 algorithm when solving this motion planning problem. The simulation result in Fig. 16 demonstrates that the generated trajectory can successfully avoid both the static and dynamic obstacles. Before reaching the goal position, the generated trajectory first avoids the dynamic obstacle, which is relatively close to the start position, then navigates around the raised part of the island. Furthermore, it is worth noting that the proposed motion planning algorithm can generate a safe, smooth and relatively short path within a very short period of time according to the summary of the simulation result provided in Table 10. Specifically, the total execution time of the main steps in the used motion planning algorithm in this simulation is less than 50 [ms] in 500 $\times$ 500 [m2] map.
In summary, the proposed D-GPMP2 algorithm can handle static and dynamic obstacles simultaneously. This characteristic significantly improves the efficiency of the proposed D-GPMP2 algorithm and broadens the potential application scenarios.
6 Implementation in ROS
This section demonstrates the practical performance of the proposed D-GPMP2 algorithm in a dynamic obstacle avoidance mission in a high-fidelity maritime environment in ROS.
6.1 Implementation details
A widely-used catamaran, the WAM-V 20 USV, is provided by the high-fidelity maritime environment and this environment is provided by [27]. Fig. 17 provides an intuitive overview of the WAM-V 20 USV above and below sea level in the ROS Gazebo virtual world. Two WAM-V 20 USVs were built inside it for a dynamic obstacle avoidance mission. As demonstrated in Fig. 18, the proposed D-GPMP2 algorithm was run on MATLAB and connected with the WAM-V 20 USVs in ROS Gazebo through ROS node.
Based on the aforementioned benchmark simulations, the proposed D-GPMP2 algorithm is able to generate an optimised path to avoid different types of dynamic obstacles. In reality or high-fidelity simulation environments, following a generated path to avoid a dynamic obstacle does not happen in a flash and takes time. As a result, we increase the input speed of a dynamic obstacle by adding a scaling factor. Despite that following a generated path consumes time, this would help guide our WAM-V 20 USV to avoid the dynamic obstacle.
6.2 Dynamic obstacle avoidance mission
During the dynamic obstacle avoidance mission, our WAM-V 20 USV transports a package from a start position to a goal position, while another WAM-V 20 USV moves towards our WAM-V 20 USV with a constant velocity. We use the proposed D-GPMP2 algorithm to generate an optimised path based on the real-time status information of the dynamic obstacle when transporting the package. By following the optimised path generated by the proposed D-GPMP2 algorithm, our WAM-V 20 USV successfully avoids the dynamic obstacle during the transportation period. In this mission, we assume the USVs have ”perfect” autopilots, meaning they follow the designated paths perfectly.
To provide a more intuitive understanding of the above process, the storyboards of the dynamic obstacle avoidance mission from the third-person perspectives is demonstrated in Fig. 19.
7 Conclusion
The work in this article successfully extended the traditional GP-based motion planning algorithm into complex and dynamic maritime environments. To achieve this purpose, a novel motion planning algorithm is proposed in this article, named as the D-GPMP2. By introducing 1) dynamic obstacle area modelling and 2) dynamic obstacle velocity modelling, we were able to use the proposed D-GPMP2 algorithm to generate safe, smooth and relatively short trajectories for USVs to avoid both static and dynamic obstacles in maritime environments in an efficient manner. Moreover, a variety of simulation environments across different platforms have been used to test the practicality and functionality of the proposed D-GPMP2 algorithm.
7.1 Future research
In terms of future research, proposed areas of focus are summarised as below:
  1. 1.
    Developing the proposed D-GPMP2 algorithm further to make it usable for AUVs in 3-D environments.
  2. 2.
    Testing the proposed D-GPMP2 algorithm on a real physical USV to further validate its practicality.